A random sample from normal population yielded sample mean=40.8 and sample standard deviation= 6.1, n = 15. H0: μ = 32.6, Ha: μ ≠ 32.6, α = 0.05. Perform the hypothesis test and draw your conclusion.

A. Test statistic: t = 5.21. P-value=0.00013. Reject H0. There is sufficient evidence to support the claim that the mean is different from 32.6.

B. Test statistic: t = 5.21. P-value=1.9E-7 (0.00000019). Reject H0. There is sufficient evidence to support the claim that the mean is different from 32.6.

C. Test statistic: t = 5.21. P-value=0.00013. Do not reject H0. There is not sufficient evidence to support the claim that the mean is different from 32.6.

D. Test statistic: t = 5.21. P-value=1.9E-7 (0.00000019). Do not reject H0. There is not sufficient evidence to support the claim that the mean is different from 32.6.
Answer #1

A random sample from normal population yielded sample mean=40.8 and sample standard deviation= 6.1, n = 15. H0: μ = 32.6, Ha: μ ≠ 32.6, α = 0.05. Perform the hypothesis test and draw your conclusion.

sample size n < 30 and σ is unknown. Hence using t distribution

Test statistic = (mean-μ)/(sample standard deviation/√n) = (40.8-32.6)/(6.1/√15) = 5.21

P value = 2*P(t>5.21) = 0.00013

A. Test statistic: t = 5.21. P-value=0.00013. Reject H0. There is sufficient evidence to support the claim that the mean is different from 32.6